In an R-L-C series circuit, the magnitude of the phase angle is 40.0o, with the source voltage behind the current. The reactance of the capacitor is 300 ohms, and the resistor resistance is 150 ohms. The average power delivered by the source is 120 W.

a. What is the reactance of the inductor?

b. What is the impedance of the circuit?

c. What is the rms current in the circuit?

d. What is rms voltage of the source?
Answer #1

CATE PAGE a Hese, x - 300w Isor R = X-X taol-40°) R + Rtao(-40%) C 14:35 Inductive veactance = 14.13.50 Impedance, 2 2 = 1958

CATE PAGE Poweo, P-12ow P= I x R 2 Р Iams R 120W isou 0.8A Ioms 0.8944. A d) P- XI oms x cosa BOS dom 125 139 v

a)Inductive reactance = 174.135 ohm

b) Impedance = 195.811 ohm

c) I​​​​​​rms = 0.8944 A

d) V​​​​​​rms = 175.139 V